OpenStudy (anonymous):

OK, last one... (z-8)/z-(9z-43)/11z

OpenStudy (anonymous):

\[\left(\begin{matrix}z-8 \\ z\end{matrix}\right)-\left(\begin{matrix}9z-43 \\ 11z\end{matrix}\right)\]

OpenStudy (anonymous):

I know the bottom number in the finished solution is 11z

OpenStudy (anonymous):

No one huh, whatever, this was the last one I needed help with, I've been stuck on it, kept skipping it, got the rest of the assignment done (aside from that last one I posted) and this is the only one left haunting me because I can't seem to figure it right.